OpenStudy (anonymous):

In triangle ABC, c = 8, b = 6, and ∠C = 60°. sin∠B = _____

OpenStudy (mathstudent55):

Use the law of sines. \(\dfrac{b}{\sin B} = \dfrac{c}{\sin C} \)

OpenStudy (anonymous):

57.4823 =57.4823?? @mathstudent55

OpenStudy (anonymous):

my answer has to be one of these how would I make it to that? @mathstudent55

OpenStudy (johnweldon1993):

So setting up the law of sines we have \[\large \frac{b}{sin(B)} = \frac{c}{sin(C)}\] We plug in what we know \[\large \frac{6}{sin(B)} = \frac{8}{sin(60^\circ)}\] And we want to solve for sin(B)

OpenStudy (johnweldon1993):

So first, we can replace sin(60) with what it equals...\(\large sin(60) = \frac{\sqrt{3}}{2}\) So \[\large \frac{6}{sin(B)} = \frac{8}{\frac{\sqrt{3}}{2}}\]

OpenStudy (johnweldon1993):

Now just cross multiply and solve for sin(B)

OpenStudy (mathstudent55):

The question asks what the sine of angle B is, not what the measure of angle B is.